Solve the following quadratic equation by completing the square:
x^2 – 6x – 3 = 0

Answer by ankor@dixie-net.com(15624)

Solve the following quadratic equation by completing the square:
x^2 – 6x – 3 = 0
:
Here it is, step-by-step, so after this, you can do it by yourself.
:
Add 3 to both sides and you have:
x^2 - 6x = +6
:
Write it, leaving a space for the value to make it a perfect square:
x^2 - 6x + __ = + 3
:
Determine this value: divide the coefficient of x (which is 6) by 2 and square it. 6/2 = 3, square 3 and you have 9, put 9 in the space. Since you added 9 on the left you have to add 9 on the right to preserve equality.
:
x^2 - 6x + 9 = 3 + 9
:
(x - 3)^2 = 12; FOIL (x-3)(x-3) to prove it to yourself
:
Find the square root of both sides and you have:
x - 3 = +/-SqRt(12)
:
x -3 = +/-SqRt(2*2*3) shows we can extract 2*2 from the radical and have:
:
x - 3 = +/- 2*SqRt(3)
:
Add 3 to both sides, we have two solutions:
x = 3 + 2*SqRt(3)
:
x = 3 - 2*SqRt(3)
:
Did this make sense to you? Any questions?